OpenStudy (anonymous):

Double check my work please? Sofia joined a golf club two years ago. She pays an annual membership fee of $1,045 and a greens fee of $35 each time she plays a game of golf. The function below can be used to calculate the total yearly golfing fee, f (g), in dollars. f (g) = 1045 + 35g, where g represents the number of times she played golf during the year. Last year she paid $1,710 as a total golfing fee. For how many games did she pay a greens fee? ----------------------------- I ended up with an answer of g=19... is that right?

OpenStudy (anonymous):

1045 + 35g = 1710 35g = 1710 - 1045 35g = 665 g = 665/35 g = 19 I agree with your answer :)

OpenStudy (anonymous):

awesome! Thank you. Could you double check another answer?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

\[\frac{ 18a ^{2}b+12ab ^{3}-24ab }{ 6ab }\]

OpenStudy (anonymous):

we were asked to find the quotient and I got 3a+2b^2-4

OpenStudy (anonymous):

3a + 2b - 4

OpenStudy (anonymous):

Thank you so much! I really appreciate the help, I'm doing an exam review haha. Thank you again!

OpenStudy (anonymous):

no problem :)